# GaussianRoom: Improving 3D Gaussian Splatting with SDF Guidance and Monocular Cues for Indoor Scene Reconstruction

### v1

@@ Abstract
Recently, 3D Gaussian Splatting(3DGS) has revolutionized neural rendering with its high-quality rendering and real-time speed. However, when it comes to indoor scenes with a significant number of textureless areas, 3DGS yields incomplete and noisy reconstruction results due to the poor initialization of the point cloud and under-constrained optimization. Inspired by the continuity of signed distance field (SDF), which naturally has advantages in modeling surfaces, we present a unified optimizing framework integrating neural SDF with 3DGS. This framework incorporates a learnable neural SDF field to guide the densification and pruning of Gaussians, enabling Gaussians to accurately model scenes even with poor initialized point clouds. At the same time, the geometry represented by Gaussians improves the efficiency of the SDF field by piloting its point sampling. Additionally, we regularize the optimization with normal and edge priors to eliminate geometry ambiguity in textureless areas and improve the details. Extensive experiments in ScanNet and ScanNet++ show that our method achieves state-of-the-art performance in both surface reconstruction and novel view synthesis.
*Equal contribution.Corresponding author.
1 Introduction
3D reconstruction from multi-view RGB images is a fundamental task in computer vision and computer graphics. The reconstructed models serve various applications such as virtual reality, video games, autonomous driving, and robotics. One typical scenario for 3D reconstruction is an indoor scene, which is characterized by large textureless areas. MVS-based methods [43, 35, 55] often produce incomplete or geometrically incorrect reconstruction results, primarily due to the geometric ambiguities introduced by the textureless areas.
Recently, neural-radiance-filed-based methods [52, 51, 62] that model the scenes using signed distance field(SDF) have achieved complete and accurate mesh reconstruction in indoor scenes, benefiting from the continuity of neural SDFs and the introduction of monocular geometric priors [62]. However, they suffer from long optimization times due to dense ray sampling in volume rendering. Fortunately, 3D Gaussian Splatting (3DGS) [24] accelerates the optimization and rendering speed of neural rendering by its differentiable rasterization technique, which also provides a new possibility in 3D scene reconstruction. Despite its impressive rendering efficiency, 3DGS often gets noisy, and incomplete reconstruction results in indoor scenes. This is primarily due to the poor initialization of the SfM point cloud in textureless regions and the under-constrained densification and optimization of Gaussians.
@@ 1 Introduction
First, we propose a SDF-guided primitive distribution strategy, which utilizes surfaces represented by SDF to guide the densification and pruning of Gaussian. For surface areas lacking initial Gaussians, we deploy new Gaussians using the SDF-guided Global Densification strategy. Simultaneously, we perform SDF-guided Densification and Pruning on existing Gaussians based on their positions relative to the scene surface. At the same time, we leverage the rendered depth from 3DGS to guide point sampling along the rays of the neural SDF, which reduces invalid sampling in free space and further improves the efficiency of optimization. Considering photometric consistency is insufficient to constrain the texture-less areas, we additionally introduce monocular normal priors in both Gaussian and neural SDF fields to regularize the geometry of textureless areas. Furthermore, to improve the details of indoor scenes, such as objects with fine structures, we propose to enhance the focus on these challenging regions during training by incorporating edge priors.
@@
  • We design an edge-aware regularization term to improve the details of reconstruction and rendering and further incorporate monocular normal priors in optimization to provide the geometric cues for textureless regions.
@@
2 Related Works
@@
Multi-view Stereo Feature-based Multi-View Stereo (MVS) methods [2, 16, 43, 45] construct explicit representation of objects and scenes by matching image features across multiple views to estimate the 3D coordinates of pixels. Surface is then obtained by applying Poisson surface reconstruction [23]. In indoor scenes, especially in large texture-less areas, these methods often struggle due to the sparsity of features. Voxel-based approaches [6, 11, 44, 33] avoid the issues of poor feature matching by optimizing spatial occupancy and color within a voxel grid, but they are limited by memory usage at high resolutions, resulting in lower reconstruction quality. Learning-based multi-view stereo methods replace feature matching [35, 49, 63] and similar processes [41] in feature-based methods with neural networks to directly predict depth or three-dimensional volumes from images [22, 55, 56, 61, 65]. However, even with the use of large data during training, errors may occur in results when dealing with occlusions, complex lighting, or areas with subtle textures.
@@
Neural Radiance Field Neural Radiance Fields (NeRF) [36] represents a scene as a continuous volumetric function of density and color, using neural networks to enable realistic new view synthesis. Methods such as Mip-NeRF [3, 54, 4] enhance rendering efficiency through optimized ray sampling techniques. Other works [37, 32, 14, 7, 29] accelerate training and rendering by leveraging spatial data structures, using alternative encodings, or resizing MLPs. Some works focus on improving the rendering quality through regularization terms. Depth regularization [12, 53], for instance, explicitly supervises ray termination to reduce unnecessary sampling time. Other approaches explore imposing smoothness constraints on rendered depth maps [38] or employing multi-view consistency regularization in sparse view settings [50, 25].
Although NeRF can produce realistic renderings for new viewpoint synthesis, the geometric quality of results directly obtained through Marching Cubes [34] is poor. Consequently, some research considers using alternative implicit functions, such as occupancy grids [39, 40] and signed distance functions (SDFs) [58, 51, 52, 57, 29], replacing NeRF’s volumetric density field. To further improve reconstruction quality,  [15, 64] propose to regularize optimization using SfM points, and  [19, 62] leverage priors such as Manhattan world assumption and pseudo depth supervision. However, these methods tend to cause missing reconstruction while consuming a long time for optimization.
@@
3D Gaussian Splatting 3D Gaussian Splatting [24] has recently become very popular in the field of neural rendering, providing an explicit representation of scenes and enabling novel view synthesis without the dependency on neural networks. During training, 3DGS consumes a significant amount of GPU memory. These efforts [13, 26] aim to compress the memory footprint of Gaussian operations. GaussianPro [9] introduces a novel Gaussian propagation strategy that guides the densification process, resulting in more compact and precise Gaussians, especially in areas with limited texture details, and similar works, such as [5, 31, 28, 20], enhance the performance of Gaussian rendering across various scenarios. DN-Splatter [48] uses depth and normal priors during the optimization process to significantly enhance reconstruction results, achieving smoother and more geometrically accurate reconstructions. Similarly, SuGaR [18], 2DGS [21], and NeuSG [8] employ Gaussian-based methods for reconstruction purposes. Our concurrent work, GSDF [60], also jointly optimizes neural SDF and 3DGS. In contrast, we focus on indoor scenes and address the challenges posed by large textureless areas in 3DGS.
3 Preliminary
3.1 3D Gaussian Splatting
@@ 3 Preliminary
3.2 Neural Implicit SDFs
NeRF [39] implicitly learns a 3D scene as a continuous volume density and radiance field from multi-view images. The method faces challenges in defining clear surfaces, often resulting in noisy surfaces when derived from density. In contrast, SDF offers a common way for representing geometry surfaces implicitly as a zero-level set, $\{\boldsymbol{x}\in\mathbb{R}^{3}\mid f_{g}(\boldsymbol{x})=0\}$, where $f_{g}(\boldsymbol{x})$ is the SDF value from an MLP $f_{g}(\cdot)$. Following NeuS [52], we replace the volume density with SDF
and convert the SDF value to the opacity $\alpha_{i}$ with a logistic function:
@@ 3.2 Neural Implicit SDFs
where $T_{i}$ is the cumulative transmittance and $N$ is the number of sample points along the ray $\boldsymbol{r}$. Similarly, the normal can be rendered as $\hat{\mathcal{N}}(p)=\sum_{i=1}^{N}T_{i}\alpha_{i}\hat{\boldsymbol{n}}_{i}$, where $\hat{\boldsymbol{n}}_{i}=\nabla f_{n}(t_{i})$ denotes the derivative of SDF at point $t_{i}$, which can be calculated by PyTorch’s automatic derivation.
4 Methodology
@@
Given multi-view posed images and point clouds obtained from the Structure from Motion (SfM) algorithm, our objective is to enable 3DGS to accurately reconstruct indoor scene geometry while preserving its rendering quality and efficiency. To this end, we first incorporate an implicit SDF field within 3DGS and design a mutual learning strategy to realize high-quality reconstruction and rendering (Sec. 4.1). Furthermore, we present the monocular geometric cues, i.e. normal prior and edge prior, to improve the reconstruction of the planes and fine details in indoor scenes respectively (Sec. 4.2). Finally, we discuss the loss functions and the overall optimization process (Sec. 4.3). An overview of our framework is provided in Fig. 1.
4.1 Mutual Learning of 3D Gaussian and Neural SDF
@@
In this section, we propose to model an indoor scene using 3D Gaussians and neural SDF with a mutual learning strategy. For 3DGS optimization, we utilize SDF to guide the distribution of Gaussian primitives, which densifies the Gaussians around the surface and significantly reduces the floaters in non-surface space. Furthermore, we introduce a Gaussian-guided sampling methodology to pilot point sampling for neural SDF, which improves the training efficiency.
@@
Figure 2: (a) Gaussian primitives distribution (b) Ground truth scene surface and Gaussian primitives distribution (c) The red Gaussian points represent new Gaussians generated by the SDF-guided Global Densification strategy, while the green Gaussian points indicate those adjusted through the SDF-guided Densification and Pruning process.
@@
SDF-guided Primitive Distribution for 3D Gaussian As illustrated in Fig. 2 (a), due to the lack of underlying geometry constrain, Gaussian primitives become disorganized during the optimization in indoor scenes, resulting in randomly scattered floaters in the non-surface areas.
SuGaR [18] introduces a regularization term to align flattened Gaussian primitives with the scene’s surface, improving mesh extraction performance but resulting in degraded rendering quality. In contrast, our approach preserves the original shape of the Gaussian primitives while utilizing the scene surface geometry information provided by the neural SDF to guide their distribution, as shown in Fig. 2 (c). Specifically, the guidance of the SDF includes two aspects, one is to deploy Gaussian primitives to achieve global densification in spatial locations with low SDF values that lack Gaussian primitives, and the other is to guide the densification and pruning of existing Gaussian primitives.
SDF-guided Global Densification. Note that the original 3DGS algorithm globally prunes Gaussian primitives based on opacity thresholds, yet it restricts its densification strategy solely to local operations like cloning or splitting, thereby posing challenges when generating new Gaussian primitives during densification in areas lacking initial Gaussians. Especially in indoor scene reconstruction tasks, such scarcity of initial Gaussians is prevalent in texture-less areas. To address this issue, we develop a global densification strategy that leverages the geometric information of the entire scene provided by the neural SDF.
@@
As depicted by the Global Densification strategy in Fig. 2 (c), we partition the scene space into $\mathrm{N}^{3}$ cubic grids and calculate the SDF value at the center of each grid. If the value falls below the threshold ($S_{c}<\tau_{s}$), it indicates that the grid is in proximity to the scene surface. Subsequently, we enumerate the existing Gaussian primitives within each grid. In cases where the number of Gaussian primitives is insufficient ($N_{g}<\tau_{n}$), we select the $K$ Gaussian neighbors of the grid’s center point and generate $K$ new Gaussian primitives within the grid. The initial attributes of these newly generated Gaussian primitives are sampled from a normal distribution defined by the mean and variance of the $K$ neighboring Gaussians.
@@
4.2 Monocular Cues constrained Optimization
It is important to note that indoor scenes encompass not only large texture-less areas such as walls and floors but also detailed regions. However, in the original 3DGS and neural SDF, the optimization only relies on image reconstruction loss without incorporating any geometric constraints, resulting in noisy and blurry results. In this section, we incorporate monocular geometric cues into our designed mutual learning pipeline, including edge prior and normal prior, which respectively are used to constrain details and flat areas.
@@ 4.3 Loss Functions
where $C_{k}$ and $\hat{C}_{k}$ are ground truth and rendered colors respectively, $w_{k}$ is the weighted term from the edge map mentioned in Sec. 4.2. The normal loss $\mathcal{L}_{normal}$ is defined as:
@@
The hyper-parameter settings are detailed in the supplementary material.
@@
5 Experiment
5.1 Experimental Setup
Dataset We evaluate the performance of our approach on reconstruction and rendering quality using 10 real-world indoor scenes from publicly available datasets: 8 available scenes from ScanNet(V2) [10] and 2 scenes from ScanNet++ [59].
Implementation Details We build our code based on 3DGS [24] and NeuS [52]. During the optimization, we divide the training into three stages. We first pre-train 3DGS with 15k iterations and then co-optimize 80k iterations, of which the 3DGS and the neural SDF don’t instruct each other for the first 6k to ensure that they each learn rough information about the scene before mutual learning. All experiments are conducted on a single Tesla V100 GPU.

### v2

@@ Abstract
Embodied intelligence requires precise reconstruction and rendering to simulate large-scale real-world data. Although 3D Gaussian Splatting (3DGS) has recently demonstrated high-quality results with real-time performance, it still faces challenges in indoor scenes with large, textureless regions, resulting in incomplete and noisy reconstructions due to poor point cloud initialization and underconstrained optimization. Inspired by the continuity of signed distance field (SDF), which naturally has advantages in modeling surfaces, we propose a unified optimization framework that integrates neural signed distance fields (SDFs) with 3DGS for accurate geometry reconstruction and real-time rendering. This framework incorporates a neural SDF field to guide the densification and pruning of Gaussians, enabling Gaussians to model scenes accurately even with poor initialized point clouds. Simultaneously, the geometry represented by Gaussians improves the efficiency of the SDF field by piloting its point sampling. Additionally, we introduce two regularization terms based on normal and edge priors to resolve geometric ambiguities in textureless areas and enhance detail accuracy. Extensive experiments in ScanNet and ScanNet++ show that our method achieves state-of-the-art performance in both surface reconstruction and novel view synthesis. Project page: https://xhd0612.github.io/GaussianRoom.github.io/
I INTRODUCTION
3D reconstruction transforms 2D images into detailed 3D models, enabling high-fidelity rendering of real-world scenes. It plays crucial roles in visual simulation for embodied intelligence, bridging the gap between simulated environments and real-world applications [27, 19, 5, 43].
One typical scenario for reconstruction is indoor scenes, which is characterized by large textureless areas. MVS-based methods [40, 49] often produce incomplete or geometrically incorrect reconstruction results, primarily due to the geometric ambiguities introduced by the textureless areas. Recently, neural-radiance-filed-based methods [48, 47, 52] that model the scenes using signed distance field (SDF) have achieved complete and accurate mesh reconstruction in indoor scenes, benefiting from the continuity of neural SDFs and the introduction of monocular geometric priors [52]. However, they suffer from long optimization times due to dense ray sampling in volume rendering. Fortunately, 3D Gaussian Splatting (3DGS) [22] accelerates the optimization and rendering speed of neural rendering by its differentiable rasterization technique, which also provides a new possibility in 3D scene reconstruction. Despite its impressive rendering efficiency, 3DGS often gets noisy, and incomplete reconstruction results in indoor scenes. This is primarily due to the poor initialization of the SfM point cloud in textureless regions and the under-constrained densification and optimization of Gaussians.
@@ I INTRODUCTION
First, we propose a SDF-guided primitive distribution strategy, which utilizes surfaces represented by SDF to guide the densification and pruning of Gaussian. For surface areas lacking initial Gaussians, we deploy new Gaussians using the SDF-guided Global Densification strategy. For the existing Gaussians, we perform SDF-guided Densification and Pruning based on their positions relative to the scene surface. At the same time, the Gaussians are used to guide point sampling of the neural SDF along the rays, which reduces invalid sampling in free space and further improves the efficiency of optimization. Secondly, considering that rendering loss is insufficient to constrain textureless areas, we additionally introduce monocular normal priors in both Gaussian and neural SDF fields to regularize the geometry of textureless areas. Furthermore, to improve the details of indoor scenes, we increase the sampling of pixels covering fine structures for training based on the edge prior.
@@
  • We design an edge-aware regularization term to improve the details of reconstruction, and further incorporate monocular normal priors in optimization to provide the geometric cues for textureless regions.
@@
II RELATED WORKS
II-A Multi-view stereo
@@
Feature-based Multi-View Stereo (MVS) methods [2, 40] construct explicit representation of objects and scenes by matching image features across multiple views to estimate the 3D coordinates of pixels. Surface is then obtained by applying Poisson surface reconstruction [21]. In indoor scenes, especially in large texture-less areas, these methods often struggle due to the sparsity of features. Voxel-based approaches [8] avoid the issues of poor feature matching by optimizing spatial occupancy and color within a voxel grid, but they are limited by memory usage at high resolutions, resulting in lower reconstruction quality. Learning-based multi-view stereo methods replace feature matching [32, 45] and similar processes [38] in feature-based methods with neural networks to directly predict depth or three-dimensional volumes from images [18, 49]. However, even with the use of large data during training, errors may occur in results when dealing with occlusions, complex lighting, or areas with subtle textures.
II-B Neural Radiance Field
@@
Neural Radiance Fields (NeRF) [33] represents a scene as a continuous volumetric function of density and color, using neural networks to enable realistic new view synthesis. Methods such as Mip-NeRF [3] enhance rendering efficiency through optimized ray sampling techniques. Other works [34, 11, 28] accelerate training and rendering by leveraging spatial data structures, using alternative encodings, or resizing MLPs. Some works focus on improving the rendering quality through regularization terms. Depth regularization [9], for instance, explicitly supervises ray termination to reduce unnecessary sampling time. Other approaches explore imposing smoothness constraints on rendered depth maps [35] or employing multi-view consistency regularization in sparse view settings [46, 23]. Although NeRF can produce realistic renderings for new viewpoint synthesis, the geometric quality of results directly obtained through Marching Cubes [31] is poor. Consequently, some research considers using alternative implicit functions, such as occupancy grids [36, 37] and signed distance functions (SDFs) [47, 48, 28, 26], replacing NeRF’s volumetric density field. To further improve reconstruction quality, [12, 53] propose to regularize optimization using SfM points, and [15, 52] leverage priors such as Manhattan world assumption and pseudo depth supervision. However, these methods tend to cause missing reconstruction while consuming a long time for optimization.
II-C 3D Gaussian Splatting
@@
3D Gaussian Splatting [22] has recently become very popular in the field of neural rendering, providing an explicit representation of scenes and enabling novel view synthesis without the dependency on neural networks. During training, 3DGS consumes a significant amount of GPU memory. These efforts [10, 24] aim to compress the memory footprint of Gaussian operations. GaussianPro [6] introduces a novel Gaussian propagation strategy that guides the densification process, resulting in more compact and precise Gaussians, especially in areas with limited texture details, and similar works, such as [4, 30, 16], enhance the performance of Gaussian rendering across various scenarios. DN-Splatter [44] uses depth and normal priors during the optimization process to enhance reconstruction results, achieving smoother and more geometrically accurate reconstructions. Similarly, SuGaR [14] and 2DGS [17] employ Gaussian-based methods for reconstruction purposes. Our concurrent work, GSDF [51], also jointly optimizes neural SDF and 3DGS. In contrast, we focus on indoor scenes and address the challenges posed by large textureless areas in 3DGS.
Figure 2: (a) Gaussian primitives distribution (b) Ground truth scene surface and Gaussian primitives distribution
III PRELIMINARY
III-A 3D Gaussian Splatting
@@ III PRELIMINARY
III-B Neural Implicit SDFs
NeRF [36] implicitly learns a 3D scene as a continuous volume density and radiance field from multi-view images. The method faces challenges in defining clear surfaces, often resulting in noisy surfaces when derived from density. In contrast, SDF offers a common way for representing geometry surfaces implicitly as a zero-level set, $\{\boldsymbol{x}\in\mathbb{R}^{3}\mid f_{g}(\boldsymbol{x})=0\}$, where $f_{g}(\boldsymbol{x})$ is the SDF value from an MLP $f_{g}(\cdot)$. Following NeuS [48], we replace the volume density with SDF and convert the SDF value to the opacity $\alpha_{i}$ with a logistic function:
@@ III-B Neural Implicit SDFs
where $T_{i}$ is the cumulative transmittance and $N$ is the number of sample points along the ray $\boldsymbol{r}$. Similarly, the normal can be rendered as $\hat{\mathcal{N}}(\boldsymbol{p})=\sum_{i=1}^{N}T_{i}\alpha_{i}\hat{\boldsymbol{n}}_{i}$, where $\hat{\boldsymbol{n}}_{i}=\nabla f_{n}(t_{i})$ denotes the derivative of SDF at point $t_{i}$, which can be calculated by PyTorch’s automatic derivation.
IV METHODOLOGY
@@
Given multi-view posed images and point clouds obtained from the Structure from Motion (SfM) algorithm, our objective is to enable 3DGS to reconstruct indoor scene geometry accurately while preserving its rendering quality and efficiency. To this end, we first incorporate an implicit SDF field within 3DGS and design a mutual learning strategy to realize high-quality reconstruction and rendering (Sec. IV-A). Furthermore, we present the monocular geometric cues, i.e. normal and edge prior, to improve the reconstruction of the planes and fine details in indoor scenes respectively (Sec. IV-B). Finally, we discuss the loss functions and the overall optimization process (Sec. IV-C). An overview of our framework is provided in Fig. 1.
IV-A Mutual Learning of 3D Gaussian and Neural SDF
@@
In this section, we propose to model an indoor scene using 3D Gaussians and neural SDF with a mutual learning strategy. For 3DGS optimization, we utilize SDF to guide the distribution of Gaussian primitives, which densifies the Gaussians around the surface and significantly reduces the floaters in non-surface space. Meanwhile, we introduce a Gaussian-guided sampling methodology to pilot point sampling for neural SDF, which improves the training efficiency.
@@
Figure 3: The red Gaussian points represent new Gaussians generated by the SDF-guided Global Densification strategy, while the green Gaussian points indicate those adjusted through the SDF-guided Densification and Pruning process.
@@
SDF-guided Primitive Distribution for 3D Gaussian As illustrated in Fig. 2, due to the lack of underlying geometry constrain, Gaussian primitives become disorganized during the optimization in indoor scenes, resulting in randomly scattered floaters in the non-surface areas. Specifically, the guidance of the SDF includes two aspects, one is to deploy Gaussian primitives to achieve global densification in spatial locations with low SDF values that lack Gaussian primitives, and the other is to guide the densification and pruning of existing Gaussian primitives.
SDF-guided Global Densification. Since the original 3DGS method densifies Gaussians based on existing ones, it struggles to generate new Gaussian primitives in regions lacking initial Gaussians. This issue becomes more pronounced in indoor scenes, where texture-less areas often exhibit a scarcity of Gaussians. To address this issue, we develop a global densification strategy that takes advantage of the geometric information of the entire scene provided by the neural SDF.
@@
As depicted by the Global Densification strategy in Fig. 3, we partition the scene space into $\mathrm{N}^{3}$ cubic grids and calculate the SDF value at the center of each grid. If the value falls below the threshold ($S_{c}<\tau_{s}$), it indicates that the grid is in proximity to the scene surface. Subsequently, we enumerate the existing Gaussian primitives within each grid. In cases where the number of Gaussian primitives is insufficient ($N_{g}<\tau_{n}$), we select the $K$ Gaussian neighbors of the grid’s center point and generate $K$ new Gaussian primitives within the grid. The initial attributes of these newly generated Gaussian primitives are sampled from a normal distribution defined by the mean and variance of the $K$ neighboring Gaussians.
@@
IV-B Monocular Cues constrained Optimization
It is important to note that indoor scenes encompass not only large texture-less areas such as walls and floors but also detailed regions. However, in the original 3DGS and neural SDF, the optimization only relies on image reconstruction loss without incorporating any geometric constraints, resulting in noisy and blurry results. In this section, we incorporate monocular geometric cues into our designed mutual learning pipeline, including edge prior and normal prior, which are used to constrain details and flat areas, respectively.
@@ IV-C Loss Functions
where $C_{k}$ and $\hat{C}_{k}$ are ground truth and rendered colors respectively, $w_{k}$ is the weighted term from the edge map mentioned in Sec. IV-B. The normal loss $\mathcal{L}_{normal}$ is defined as:
@@
V Experiment
V-A Experimental Setup
Dataset We evaluate the performance of our approach on reconstruction and rendering quality using 10 real-world indoor scenes from publicly available datasets: 8 scenes from ScanNet(V2) [7] and 2 scenes from ScanNet++ [50].
Implementation Details We build our code based on 3DGS [22] and NeuS [48]. During the optimization, we divide the training into three stages. We first pre-train 3DGS with 15k iterations and then co-optimize 80k iterations, of which the 3DGS and the neural SDF don’t instruct each other for the first 6k to ensure that they each learn rough information about the scene before mutual learning. During the co-optimization stage of 3DGS and neural SDF, we implement the SDF-guided Densification and Pruning strategy every 100 iterations and the SDF-guided Global Densification strategy every 2000 iterations, rather than employing the original Gaussian densification and pruning strategy, with other hyper-parameters remaining mostly consistent with 3DGS. As for the neural SDF, We sample 1024 rays per batch and 64+64 points on each ray. After training, we extract a mesh from the SDF by the Marching Cube algorithm [31] with the volume size of $512^{3}$. For geometry comparison, we utilize Poisson reconstruction [20] to get the mesh for COLMAP [39] and some Gaussian-based methods.